OpenStudy (anonymous):

Rewrite with only sin x and cos x. cos 3x

OpenStudy (anonymous):

\[\cos 3x=\cos \left( 2x+x \right)=?\] cos(a+b)=cos a cos b-sin a sin b

OpenStudy (anonymous):

sin 2a =2sin a cos a

OpenStudy (anonymous):

I'm sorry, I'm lost

OpenStudy (anonymous):

you want to write formula for cos 3x or you want to derive the formula.

OpenStudy (anonymous):

The question asks to rewrite cos 3x

OpenStudy (anonymous):

\[\cos 3x=4\cos ^{3}x-3\cos x\] if you want to derive this formula then proceed as written above.

OpenStudy (anonymous):

@jdoe0001